OpenStudy (anonymous):

The first challenge to complete is the Penny Problem. The radio station is giving the winner of this challenge a prize pack that includes tickets to see his or her favorite band in concert. To start off the challenge, the radio station has placed pennies in a cylindrical glass jar. Each penny is 0.75 inches in diameter and 0.061 inches thick. If the cylindrical glass jar containing the pennies has a diameter of 6 inches and a height of 11.5 inches, how many pennies can fit inside the jar? You must show all work to receive credit.

OpenStudy (anonymous):

diameter = 6 in. height = 11.5 in. The formula you have chosen is correct as: Volume(cylinder) = pi * r^2 * h V = 3.14 * 3^2 * 11.5 V = 324.99 in. ^3 For the above calculation, you got 18. So, I think there is an error in the way you are multiplying the numbers in your formula. Recheck to be sure you are squaring the radius, substituting a value of 3.14 for pi, and multiplying everything together to get volume.

OpenStudy (anonymous):

thats what my teacher said